OpenStudy (anonymous):

Solve this by Completing Square a^2+6a=8 Pls show the solution clear.

OpenStudy (anonymous):

I'm sorry if i'm wrong, but you should be simplifying it. So you subtract 6a on both sides. and since the a is squared, you have to find the square root of it. so your answer should be a=(sqrt)8-6a Again, i'm sorry if i'm not very good at explaining it. Lol

OpenStudy (anonymous):

a^2 +6a =8 divide the middle term by 2........... ( 6/2 =3) then square 3^2 =9 add 9 to both sides of the equation........a^2+6a+9=17.......The left hand side is a perfect square and can be factored (a+3)^2=17 now sqrt both sides a+3=+/- sqrt(17)......move the 3 over and you are finished a=-3 +/- sqrt(17)

OpenStudy (anonymous):

it is important to put the +/- in front of anything you square root in this process, this will give you two answers, good luck.

OpenStudy (anonymous):

tyvm

OpenStudy (anonymous):

a^2+6a =8 a(a+6) =8 a=8 or a=2